OpenStudy (anonymous):

Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid one hundred thirty-five dollars. This year, he paid one hundred sixty nine dollars. What was the percent increase in the cost of his plane ticket? Show your work.

OpenStudy (anonymous):

Is anybody there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

o sorry

OpenStudy (anonymous):

its ok x/100 = $12/($9+$8+$11+$12) 1200=40*x x=30% olive oil=75% or 0.75 3 = 0.75*x x=4 (total solution volume is 4 cups) $80,000*0.15=$12,000 $12,000+$40,000=$52,000 $58,000 is more than $52,000 so Isabelle made more money than Randall. 143-130=13 13 = 0.01*x*130 13 = 1.3*x 10 = x (10% increase)

OpenStudy (anonymous):

got it from yahoo answers

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

your welcome don't forget to medal!:)

OpenStudy (anonymous):

just click the best response button (if you want)